OpenStudy (anonymous):

If i have 3 over the (sqrt) 3 would it just be 3 since 3 is prime?

OpenStudy (anonymous):

\[(3\sqrt{3} )/3\]

OpenStudy (anonymous):

3/(sqrt)3

OpenStudy (radar):

\[3 \over \sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (radar):

Simplify by rationalizing the denomintor

OpenStudy (anonymous):

i did but im not sure if my answer is correct

OpenStudy (anonymous):

is it 3

OpenStudy (anonymous):

\[{{3}\over{\sqrt{3}}}= {{\sqrt{3}\times \sqrt{3}}\over{\sqrt{3}}}=\sqrt{3}\]

OpenStudy (anonymous):

ohh okay thanks